marcoeconomis Suppose that in Zambia, people consume only two goods, peas and beans. Suppose that John, a citizen in that country has an income of ZMW10, the price of beans is 20N per kg and the price of peas is 40N per kg. a) Suppose that John consumes 30 kg of beans. Assume that he wants to spend all his income, how many kg of peas is he going to consume? b) Assume that the price of peas falls from 40N to 20N. Assuming that John still consumes 30kg of beans, find the new quantity of peas. c) After the decrease in the price of peas to 20N, assume that Jonh is just as well off as he was in (a) if he has an income of ZMW7.6. However, with that income and the new price of peas he would have consumed 20 kg of beans. Find the quantity of peas he would have consumed in this case. d) Find the substitution effect on consumption of peas due to the decrease in the price of peas in (c). e) Find the income effect on consumption of peas due to the decrease in income in (c).
1
Expert's answer
2020-09-23T09:43:43-0400

a) If John consumes 30 kg of beans. and wants to spend all his income, then he is he going to consume "Qp = (10 - 0.2\u00d730)\/0.4 = 10."

b) If the price of peas falls from 40N to 20N assuming that John still consumes 30kg of beans, then the new quantity of peas is "Qp = (10 - 0.2\u00d730)\/0.2 = 20." . c) If he has an income of ZMW7.6 and he would have consumed 20 kg of beans, then the quantity of peas he would have consumed in this case is:

"Qp = (7.6 - 0.2\u00d720)\/0.4 = 9."

d) The substitution effect on consumption of peas due to the decrease in the price of peas in (c) is

"(0.4 - 0.2)\u00d720 = 4."

e) The income effect on consumption of peas due to the decrease in income in (c) is 10 - 7.6 = 2.4.